OpenStudy (anonymous):

f(x) = 2e^3x + 1 Find the inverse f^-1(x). Can anyone help?

OpenStudy (jackellyn):

Set it equal to y. Switch the x and the y, solve for y.

OpenStudy (jackellyn):

y=2x^3+1 x=2y^3+1 Solve for y.

OpenStudy (anonymous):

thx @jackellyn

OpenStudy (jackellyn):

Did you get the answer?

OpenStudy (anonymous):

my calculator is messing up, can you walk me through it pls. @jackellyn

OpenStudy (jackellyn):

so to solve for y you need to perform reverse operations. If we have x=2e^3y +1 What do we need to do first to get y alone?

OpenStudy (anonymous):

subtract 3y on both sides @jackellyn i think

OpenStudy (jackellyn):

Well the 3y is part of the exponent so we can't get rid of it just yet. First we need to get rid of the +1 and move it to the other side so do we add it to both sides or subtract it?

OpenStudy (anonymous):

subtract it @jackellyn

OpenStudy (jackellyn):

Right so we have x-1= 2e^3y Now we need to get rid of the 2 in front of the e^3y. What do you get?

OpenStudy (anonymous):

divided

OpenStudy (anonymous):

@jakellyn if thats correct

OpenStudy (anonymous):

@jackellyn you there

OpenStudy (jackellyn):

Yes, divide so then we have (x-1)/2=e^3y Do you remember the opposite operation of e?

OpenStudy (anonymous):

I don't recall @jackellyn

OpenStudy (jackellyn):

You need to take the natural log of both sides to get rid of the e and stay only with 3y. ln (x-1/2)=ln e^3y ln (x-1/2)=3y Solve for y.

OpenStudy (anonymous):

\[y =\frac{ x }{ 3}-\frac{ 1 }{ 6 }\] is that correct? @jackellyn

OpenStudy (anonymous):

so is this the inverse for \[f ^{-1}\]

OpenStudy (anonymous):

@jackellyn idk if you saw my replies^^^^

OpenStudy (jackellyn):

Sorry, I didnt. You can't get rid of the ln. It stays all together and then you can divide the 3. So \[y=\frac{(\ln \frac{ x-1 }{ 2 }) }{ 3 }\]

OpenStudy (anonymous):

so thats the inverse of \[f ^{-1}(x)\]

OpenStudy (jackellyn):

It is the inverse of f(x). \[f ^{-1}(x)\] is the notation for inverse.

OpenStudy (anonymous):

okay i understand now thx a lot

OpenStudy (anonymous):

@jackellyn one more question what is the inverse